# A Survey on Large Language Models with Multilingualism: Recent Advances and New Frontiers

### v1

@@ 4.6 Limitations and Future Directions on Inference Strategies
Universal inference paradigm. Previous PLMs are task-specific and language-specific, indicating that it is essential to specify the language ID. For instance, we need to assign the language ID of input and output respectively when leveraging the M2M-100 for machine translation Fan et al. (2021). However, it is not required in the existing paradigm of LLMs, such as GPT-4. Due to the universality of LLMs, an aim is to keep them as task-agnostic and language-agnostic as feasible. Current LLMs employ a standardized approach for inference across all languages (either direct inference, pre-translation inference, CoT, or RAG). However, as shown in the experiments and analyses above, the inference strategies exhibit varied performances when faced with different LLMs, tasks, and languages. Therefore, a flexible and universally applicable paradigm is desirable to uniformly handle the environments of all languages (i.e., high-/medium-/low-resource languages).
@@
Emergence ability. One important advantage of LLMs is the emergence ability that is informally defined as “the ability that does not exist in small models but appears in large models” Wei et al. (2022b). However, this advantage is not embodied significantly in multilingual scenarios. Thus, the performance of LLMs is far behind that of the English scenario, and difficult to outperform small language models. It is not only due to limitations during the training phase but also potentially influenced by inappropriate strategies during the inference phase.
@@
Model collaboration. Existing inference pipelines are designed for single LLMs. Due to data scarcity in low-resource languages, it is challenging to address tasks for all languages within a single model Magueresse et al. (2020); Ranathunga et al. (2023). Considering that the smaller models excel at specific tasks or handling specific languages, which do not require large-scale data to optimize Miceli-Barone et al. (2017); Dabre et al. (2019); Bapna and Firat (2019). It is beneficial to leverage the strengths of both large and small models and solve tasks for different resource languages through model collaboration. The model collaboration that effectively bridges the large and small models is a direction yet to be explored for enhancing inference performance in the multilingual scenario.
@@ 5 Security of Multilingual Large Language Models
JailBroken Wei et al. (2024) GPTFUZZER Yu et al. (2023b) AutoDAN Liu et al. (2023c) DeepInception Li et al. (2023e) ICA Wei et al. (2023b) PAIR Chao et al. (2023) ReNeLLM Ding et al. (2023)
@@
GCG Zou et al. (2023) Multilingual Deng et al. (2023) Cipher Yuan et al. (2024) CodeChameleon Lv et al. (2024)
GPT-3.5-turbo 12% 12% 100% 35%
GPT-4-0613 0% 0% 58% 0%
LLaMA-2-7B-Chat 46% 46% 6% 31%
LLaMA-2-13B-Chat 46% 46% 4% 41%
Vicuna-7B-v1.5 94% 94% 100% 93%
Vicuna-13B-v1.5 94% 94% 100% 94%
ChatGLM-3 34% 34% 95% 85%
Qwen-7B-Chat 48% 48% 100% 82%
Intern-7B 10% 10% 100% 92%
Mistral-7B 82% 82% 100% 99%
Table 4: An overview of Greedy Coordinate Gradient, Prompt-Based and Multilingual attack methods jailbreaking for LLMs on the AdvBench Zhou et al. (2024b). The evaluation method is consistent with the EasyJailbreak Zhou et al. (2024b) framework, which uses GPT-4-turbo-1106 as the scoring model and the evaluation prompts from GPTFUZZER Yu et al. (2023b).
@@ 5.1 Attack Methods
To explore the security of LLMs, a red team attack on LLMs is required. A red-team attack is a cybersecurity exercise where a group of ethical hackers, known as the red team, simulate real-world cyberattacks on an organization’s systems, networks, or infrastructure. The goal of a red-team attack is to identify vulnerabilities, weaknesses, and potential security breaches that malicious actors could exploit Mirkovic et al. (2008). A common practice is the “jailbreak” attack which typically refers to the unauthorized access or modification of the underlying code or functionality of models. Essentially, it involves breaking out of the constraints or limitations imposed by the design or usage policies of LLMs. It includes techniques to bypass security measures or enable functionalities that are not intended or permitted by the developers.
According to the criteria in existing studies Wu et al. (2024); Zhou et al. (2024b), the jailbreak methods on LLMs can be divided into three types: Greedy Coordinate Gradient (GCG) jailbreak Sitawarin et al. (2024); Zou et al. (2023), prompt-based jailbreak Wei et al. (2024); Liu et al. (2023b); Shen et al. (2023); Deng et al. (2024); Li et al. (2024b); Puttaparthi et al. (2023); Liu et al. (2023c); Jin et al. (2024) and multilingual jailbreak Shen et al. (2024); Deng et al. (2023); Puttaparthi et al. (2023); Yong et al. (2023); Xu et al. (2023); Li et al. (2024c); Yuan et al. (2024); Huang et al. (2023d). In this section, we focus on these three jailbreak methods, especially the multilingual jailbreak. The first two methods involve generic attacks on LLMs, and the latter emphasizes jailbreaking through multiple languages. All these methods are aimed at bypassing the security measures of LLMs to generate malicious information. Since most of the jailbreaking methods are customized, the effectiveness of each method varies depending on the specific LLMs. To provide a comprehensive comparison, we investigate the performance of different jailbreaking methods across various LLMs based on a unified evaluation framework, as shown in Table 4.
@@ 5.1.1 Greedy Coordinate Gradient Jailbreak
The Greedy Coordinate Gradient  Zou et al. (2023) is an attack method based on greedy algorithms and gradient metric design hints. It first creates a seed prompt, and then iteratively replaces the tokens. The best prompt for attacking the model is determined by calculating the gradient. Sitawarin et al. Sitawarin et al. (2024) propose GCG++ to increase the attack success rate (ASR) of the original GCG method on LLaMA-2-7B. On the basis of the original GCG, they replace the cross entropy loss with multi-class loss to avoid vanishing gradients in the $Softmax$, which is effective in attacking LLM. Another discovery is that LLMs such as LLaMA-2-7B are format sensitive, having a strong prior for predicting a space token at the beginning of the model response. GCG++ adds a simple format-ware target string (i.e., “[ASSISTENT]:” before the input prompt to force the model to output harmful content. Through the two changes, GCG++ increases the GCG attack success rate on LLaMA-2-7B from 56% to 80%.
@@ 5.1.2 Prompt-Based Jailbreak
The prompt-based methods aim to mislead LLM into generating harmful content based on specifically designed templates. Among them, LLM-based Liu et al. (2023c); Mehrotra et al. (2023); Sitawarin et al. (2024); Chao et al. (2023); Deng et al. (2024); Yu et al. (2023b); Ding et al. (2023); Bhardwaj and Poria (2023) and rule-based Wei et al. (2024); Li et al. (2023e); Liu et al. (2023b); Shen et al. (2023); Puttaparthi et al. (2023) are the two common used methods for designing attacking prompts. The differences between GCG, LLM-based, and rule-based are presented in Table 5 with various aspects.
Rule-based jailbreak methods are those works that design specific prompts to formulate rules for LLMs Wei et al. (2024); Li et al. (2023e), or directly collect prompts from various channels such as websites and forums Liu et al. (2023b); Shen et al. (2023); Puttaparthi et al. (2023). Rule-based jailbreak methods rely more on human participation because these rules are customized based on empirical experience in specific scenarios. Thus, the produced prompts are effective in a narrow area but lack versatility and are static, which can be easily defended.
LLM-based jailbreak methods leverage the instances of LLMs to optimize existing attack prompts or generate new jailbreak prompts. Different from the GCG series approaches, the principle of LLM-based jailbreak methods acts as using an agent LLM to attack a target LLM. Existing studies  Liu et al. (2023c); Mehrotra et al. (2023); Sitawarin et al. (2024); Chao et al. (2023) usually take the agent LLM as an optimizer, scoring module or evaluator to optimize prompts for attackers. Some other methods Deng et al. (2024); Yu et al. (2023b); Ding et al. (2023); Bhardwaj and Poria (2023) use the generation capabilities of LLM to perform multiple operations such as rewriting and shortening feasible attack prompts. They also wrap harmful prompts with scenarios to avoid safe alignment. These operations based on LLMs can enable the original attack prompt to evade the security alignment policy.
@@ 5.1.3 Multilingual Jailbreak
Due to the unequal security alignment of LLMs across different languages, with higher security in high-resource languages compared to low-resource languages, the multilingual jailbreak methods Shen et al. (2024); Deng et al. (2023); Yong et al. (2023); Xu et al. (2023); Li et al. (2024c) attend to cheat the LLMs via the alignment vulnerabilities between high-resource and low-resource language. Some other methods Yuan et al. (2024); Huang et al. (2023d); Lv et al. (2024); Wei et al. (2024) use special encoding or ciphers to disguise harmful prompts to achieve jailbreak, which can be considered as a type of special language.
@@
The existing studies on the potential security risks of LLMs, primarily focused on the usage of English scenarios. Under the multilingual scenario, Deng et al. Deng et al. (2023) reveal the presence of multilingual jailbreak challenges within LLMs, who divide potential risks into two scenarios: (1) unintentional and (2) intentional. The unintentional scenario involves users inadvertently bypassing the safety alignment by querying LLMs with non-English prompts, while the intentional scenario refers to users intentionally combining malicious instructions with multilingual prompts. Compared to high-resource languages, low-resource languages are approximately three times more likely to encounter harmful content in GPT-3.5 and GPT-4, and thus the LLMs are much easier to attack via low-resource languages. This could also be the reason why the model attack approaches are designed with low-resource languages within these two scenarios.
@@
For the unintentional scenario, Shen et al. Shen et al. (2024) indicate that the bottleneck in cross-lingual alignment lies within the training stage of LLMs. Thus, they study the effect of instruction tuning with RLHF and supervised fine-tuning (SFT) on the HH-RLHF dataset Bai et al. (2022). The results show that although training with high-resource languages can improve model alignment, the effect of training with low-resource languages is still negligible. For the intentional scenario, various prompts are designed to bypass the security defenses of the LLMs via high-resource languages. In particular, Yong et al. Yong et al. (2023) successfully circumvent the safeguard of GPT-4 by translating unsafe English prompts into low-resource languages. They propose the LRL-Combined Attacks approach to achieve 79% ASR on the AdvBench dataset Zou et al. (2023). Xu et al. Xu et al. (2023) study a new type of black box jailbreak attack, Cognitive Overload, which is specifically designed for the cognitive structure and processes of LLMs. With Google Cloud API, they translate original English harmful instructions from AdvBench and MasterKey Deng et al. (2024) into 52 other languages and propose multilingual cognitive overload Sweller (1988, 2011) to hinder in-context learning and inference process when the knowledge exceeds the limited capacity of LLMs Szulewski et al. (2021). By using Google Cloud API to translate English to other 52 languages and nllb-200-distilled-1.3B to translate other non-English responses back to English, they conducted a multilingual version of their attack method. With multilingual Cognitive Overload, their results show that LLMs are more vulnerable to non-English adversarial prompts. The observation indicates that the further a language is from English, the more effective the malicious prompt conveyed is in attacking LLMs. Moreover, Li et al. Li et al. (2024c) conduct an extensive empirical study on multilingual jailbreak attacks, which develops a novel semantic preservation algorithm to create multilingual jailbreak prompts. With these prompts, they reveal patterns in multilingual jailbreak attacks and implement fine-tuning mitigation methods for defending against cross-lingual jailbreak attacks.
@@
Password jailbreak. From the perspective of writing scripts, special encoding or password can be considered as a language conversion, i.e., an English prompt can be converted into a password representation by encryption. Though the password representation can also be converted to English or other languages via decryption, the encrypted prompts cannot be understood directly by humans but by the LLMs, and thus could lead to secure alignment fails. For example, Huang et al. Huang et al. (2023d) propose the generation exploitation attack to manipulate variations of decoding methods (e.g., greedy decoding and sampling-based decoding), which can easily disrupt model alignment because existing alignment procedures are based on default decoding settings. Vulnerabilities are exposed when configurations of decoding change slightly. Yuan et al. Yuan et al. (2024) propose a novel framework CipherChat to examine the generalizability of safety alignment to non-natural languages, where the user can chat with LLMs through cipher prompts. Furthermore, Lv et al. Lv et al. (2024) introduce a hypothesis for the safety mechanism of LLMs: intent security recognition followed by response generation. Following this hypothesis, they propose CodeChameleon to transform queries into decryptable formats with custom Python functions. This approach enables the modification of the original query and achieves state-of-the-art ASR on 7 LLMs. Besides, Wei et al. Wei et al. (2024) propose Jailbroken, which obfuscates the queries using base64 to bypass the safety training of LLMs. They observe that LLMs can understand base64 but cannot defend against attacks under base64.
@@
Other jailbreak. Apart from the aforementioned methods, Shayegani et al. Shayegani et al. (2023) combine an image targeted towards toxic embeddings with generic prompts to accomplish the jailbreak, which utilizes four embedding space target strategies to poison the vision encoder. Their attacks achieve 87% and 63.3% ASR on two vision-language models, LLaVA Liu et al. (2024b) and LLaMA-Adapter V2 Gao et al. (2023b), respectively. Rando et al. Rando and Tramèr (2023) embed a “jailbreak backdoor” into the LLM by poisoning the RLHF training data. Then, users can easily achieve jailbreak by using a trigger word like the “sudo” command in the Linux system. Wolf et al. Wolf et al. (2023) propose a theoretical approach Behavior Expectation Bounds (BEB) that investigates several limitations of alignment in LLMs, which exposes fundamental problems and emphasizes the necessity of designing reliable mechanisms to ensure AI safety.
Model Attacker: PAIR Chao et al. (2023)
No Defense SmoothLLM Robey et al. (2023) Perplexity Filter Jain et al. (2023)
@@
Table 6: An overview of defense methods under jailbreaking on listed LLMs. The evaluation method for the Attack Success Rate (ASR) indicator is consistent with JailbreakBench Chao et al. (2024). The responses of LLMs are evaluated using LLaMAGuard-7B.

### v2

@@ 4.6 Limitations and Future Directions on Inference Strategies
Universal inference paradigm. Previous PLMs are task-specific and language-specific, indicating that it is essential to specify the language ID. For instance, we need to assign the language ID of input and output respectively when leveraging the M2M-100 for machine translation [345]. However, it is not required in the existing paradigm of LLMs, such as GPT-4. Due to the universality of LLMs, an aim is to keep them as task-agnostic and language-agnostic as feasible. Current LLMs employ a standardized approach for inference across all languages (either direct inference, pre-translation inference, CoT, or RAG). However, as shown in the experiments and analyses above, the inference strategies exhibit varied performances when faced with different LLMs, tasks, and languages. Therefore, a flexible and universally applicable paradigm is desirable to uniformly handle the environments of all languages (i.e., high-/medium-/low-resource languages).
@@
Emergence ability. One important advantage of LLMs is the emergence ability that is informally defined as “the ability that does not exist in small models but appears in large models” [346]. However, this advantage is not embodied significantly in multilingual scenarios. Thus, the performance of LLMs is far behind that of the English scenario, and difficult to outperform small language models. It is not only due to limitations during the training phase but also potentially influenced by inappropriate strategies during the inference phase.
@@
Model collaboration. Existing inference pipelines are designed for single LLMs. Due to data scarcity in low-resource languages, it is challenging to address tasks for all languages within a single model [347, 348]. Considering that the smaller models excel at specific tasks or handling specific languages, which do not require large-scale data to optimize [349, 350, 351]. It is beneficial to leverage the strengths of both large and small models and solve tasks for different resource languages through model collaboration. The model collaboration that effectively bridges the large and small models is a direction yet to be explored for enhancing inference performance in the multilingual scenario.
5 Multilingual Information Retrieval
The task of Information Retrieval (IR) is to find the relevant documents that satisfy the information needs of the users (in the form of queries) from a large collection [352]. Multilingual IR studies the IR tasks in more than only English, and is generally categorized into the monolingual, cross-lingual and multilingual scenarios based on the language(s) of the queries and documents: Given a query in the language ($L_{1}$), monolingual retrieval aims to find relevant information from the same language as the query ($L_{1}$), cross-lingual retrieval finds relevant information from a different language ($L_{2}$), and multilingual retrieval finds relevant information from multiple different languages ($L_{i},i\in\{1,...,n\}$). There are many surveys and discussions related to this section, including multilingual retrieval progress prior to neural IR [353, 354, 355, 356, 357, 358, 359, 360, 361], progress on general neural IR prior to the emergence of LLMs [362], and interplay of LLM and IR [363, 364, 365]. This section will focus on the multilingual aspect, especially the new opportunities of multilingual retrieval brought by LLMs (“LLM for mIR”), as oppose to the RAG methods (“mIR for LLM”) introduced in the Section 4.5 above.
5.1 Synthetic Training Data
Synthetic datasets for multilingual retrieval are traditionally created in two approaches: machine translation [115, 366], and natural semantic structures [367], e.g., title–passage [368], inner-connected Wikipedia links [369, 370], and so on.
Spearheaded by InPars [89, 90, 91] and Promptagator [92], LLMs bring the third approach, which generates large-scale synthetic data for training retrieval models in an affordable way. While the above two works focus on English, following this line, SwimIR [93] builds large-scale training data for both cross-lingual and monolingual retrieval tasks. JH-POLO [94] build cross-lingual retrieval training data by generating English queries based on non-English positive and negative passages using LLM. It has also been found that synthetic data generated by LLM could improve the performance of LLM-based embedding models: mE5-Mistral [95] generate synthetic data by GPT-3.5 and GPT-4 on 93 languages. Gecko [96] built Few-shot Prompted Retrieval dataset (FRet). Both works adopt a two-stage generation pipeline: in the first stage, the LLM generates the query and the task description(s) given passage(s); in the second stage, while both works produce positive and (hard) negative passages, Gecko asks LLM to score the positive and negative passages from candidates returned by a retriever, while mE5-Mistral lets the LLM generate the positive and negative passages based on given requirements. Additionally, Arctic-Embed [97] find that hard negative mined from existing corpora is in higher quality compared to the ones generated by LLM.
5.2 Multilingual Retrievers
According to the taxonomy by Lin and Ma [371], the retrievers are categorized into the unsupervised sparse, supervised sparse and, supervised dense models, where the dense models could be further categorized into the single-vector and multi-vector models. This section briefly introduces models under each category that are extended into multilingual scenarios, and the impact of LLMs on the category if applicable.
The unsupervised sparse models refers to the bag-of-words exact matching ranking algorithms, e.g., BM25 [372]. It requires language-specific analyzers to extend the models into the multilingual scenarios, which are designed based on expert knowledge on the target language. This category of models serves a strong baseline itself on the mono-lingual retrieval task [373], or on the cross-lingual and multilingual retrieval when used together with the query or document translation [374, 375, 376].
The supervised sparse models learn the weight per term from the training data rather than using the collection statistics as unsupervised sparse models [377, 378, 379, 380, 381, 382]. Works extending this category into the multilingual scenarios are mainly surrounding SPLADE on monolingual retrieval [383] or cross-lingual retrieval [384, 385]. All above works are based on BERT-level multilingual pretrained models.
The supervised dense models category, including both the single-vector and multi-vector dense retrieval models, resides most of the work on the multilingual IR and LLM-related embedding models. The dense models was firstly extended into the multilingual scenarios by switching the backbones into mBERT or XLM-R [373, 386, 387]. Many methods that are found effective on English also show similar performance improvements on the multilingual scenarios, for example, extending the pretraining and pre-finetuning corpora into multilingual unsupervised data or translated corpora [115, 387], knowledge distillation from rerankers [387] or even from English retriever [388]. Recently, more multilingual embedding models are proposed with large-scale pre-training and pre-finetuning. Open-source models include mE5 [98], mGTE [99], and BGE [100], whereas [101, 102, 103] provide black-box API for the multilingual embedding models.
Many LLM-based embedding models emerge these years: Focusing on the retrieval task, RepLLAMA [104] shows that LLM-based embedding models could be fine-tuned to achieve better in-domain effectiveness and out-of-domain generalizability. PromptReps [105] shows that the LLMs could be prompted to generate dense and sparse representations to achieve competitive zero-shot performance on the passage retrieval tasks. More works emerge with the target of unifying multiple embedding tasks into a single model [106]. The modification on the model design includes: enabling bi-directional attention [107, 108, 109], adding instructional tuning [107, 110], adapting Matryoshka representations [101, 111], and merging multiple models [112]. Training wise, AnglE [113] proposed angle-optimized text embedding models to eliminate the gradient saturation zone of cosine function. Yet most of the work along this line focuses on English embedding models, only Gecko [96] supports the multilingual inputs. Additionally, NLLB-E5 [114] propose to integrate NLLB models with E5 to enable the multilingual ability of English embedding models.
Works extending multi-vector dense retrieval models into multilingual scenarios mainly focus on ColBERT [389], dubbed mColBERT [115, 116] or ColBERT-X [117]. The extension covers monolingual retrieval [115, 122, 116], cross-lingual retrieval [117, 118, 119], and multilingual retrieval [120, 121]. Other than adopting the default model structure, ColBERT-XM [122] adds modular language-specific adapter in XMOD architecture, which shows competitive zero-shot monolingual retrieval performance. BGE-M3 [123] unites the single-vector dense, multi-vector dense and sparse models by distilling the knowledge from the ensemble of the three types of models into each single models, demonstrating strong on both monolingual and cross-lingual retrieval tasks.
5.3 Multilingual Rerankers
Early works multilingual neural rerankers are mostly based on mBERT [390], exploring the translation configuration on the training and inference stages [391, 392]. Later works found that mMiniLM and mT5 serve as better backbones as multilingual rerankers [115, 124].
Recently, a line of works explored using LLM as the zero-shot rerankers in retrieval initiated by RankGPT [125] and LRL [126], dubbed listwise rerankers, followed by works on distilling from the GPTs with enhanced training techniques [128, 129], and on building listwise reranker without relying on the close-sourced GPT-models [130]. One criticism on the listwise rerankers is the sequential inference step caused by the sliding-window aggregation step that merges the ranked documents from different batch. To address the issue, different inference and aggregation strategies have been proposed, including Setwise [127] and TourRank [131]. Limited works explored the listwise rerankers on the multilingual IR tasks, the only one to our best knowledge is Adeyemi et al. [132], which evaluated GPT-based listwise reranker on CIRAL [393], a cross-lingual retrieval benchmark on African languages, finding that GPT-4 yields competitive zero-shot performance on the task and even on par with the zero-shot results with machine-translated documents on some of the languages.
5.4 Challenges and Future Directions on Multilingual Information Retrieval
The aforementioned existing studies demonstrate how LLM-based methods enhance the effectiveness and out-of-domain generalization in information retrieval tasks. Beyond integrating LLMs as components in the established retrieval-reranking pipeline, LLMs also open new possibilities to the paradigm of search process and broader information access: For example, Li et al. [394] and Lee et al. [395] both propose a unified generative framework that integrates retrieval and question answering, where retrieval is considered as an inherent part together with the other generative tasks rather than as a separate component as in RAG.
That said, challenges present in deploying LLM in assessible search systems, which include inherently high latency for indexing and searching, as well as high computational demand during inference and fine-tuning. Current work that distills knowledge from LLM to smaller models strike a balance between effectiveness and efficiency [396].
Beyond the general challenges of LLM-based retrieval, The open-source state-of-the-art LLM usually have fewer supported languages compared to encoder-only or encoder-decoder models: LLaMA-3 [397] and Command-R+https://docs.cohere.com/v2/docs/command-r-plus support around 20 languages, while encoder-only models such as mBERT and XLM-R support 100 languages, and translation-targeting encoder-decoder such as NLLB support 200 languages. Current retrieval methods are applied to LLMs, which mainly regard LLMs as a knowledge store. However, in low-resource languages, LLMs lack generation capabilities and have not been trained with large-scale data, thus they are difficult to serve as a reliable source of knowledge. How can the above methods be applied to languages not yet supported by LLM but available in smaller language models?
@@ 6 Security of Multilingual Large Language Models
JailBroken [135] GPTFUZZER [398] AutoDAN [140] DeepInception [399] ICA [400] PAIR [401] ReNeLLM [402]
@@
GCG [134] Multilingual [143] Cipher [148] CodeChameleon [403]
GPT-3.5-turbo 12% 100% 80% 90%
GPT-4-0613 0% 63% 75% 72%
LLaMA-2-7B-Chat 46% 2% 61% 80%
LLaMA-2-13B-Chat 46% 0% 90% 67%
Vicuna-7B-v1.5 94% 94% 28% 80%
Vicuna-13B-v1.5 94% 100% 76% 73%
ChatGLM-3 34% 100% 78% 92%
Qwen-7B-Chat 48% 99% 58% 84%
Intern-7B 10% 99% 99% 71%
Mistral-7B 82% 100% 97% 95%
Table 4: An overview of Greedy Coordinate Gradient, Prompt-Based and Multilingual attack methods jailbreaking for LLMs on the AdvBench [404]. The evaluation method is consistent with the EasyJailbreak [404] framework, which uses GPT-4-turbo-1106 as the scoring model and the evaluation prompts from GPTFUZZER [398].
@@ 6.1 Attack Methods
To explore the security of LLMs, a red team attack on LLMs is required. A red-team attack is a cybersecurity exercise where a group of ethical hackers, known as the red team, simulate real-world cyberattacks on an organization’s systems, networks, or infrastructure. The goal of a red-team attack is to identify vulnerabilities, weaknesses, and potential security breaches that malicious actors could exploit [405]. A common practice is the “jailbreak” attack which typically refers to the unauthorized access or modification of the underlying code or functionality of models. Essentially, it involves breaking out of the constraints or limitations imposed by the design or usage policies of LLMs. It includes techniques to bypass security measures or enable functionalities that are not intended or permitted by the developers.
According to the criteria in existing studies [153, 404], the jailbreak methods on LLMs can be divided into three types: Greedy Coordinate Gradient (GCG) jailbreak [133, 134], prompt-based jailbreak [135, 136, 137, 138, 139, 144, 140, 141] and multilingual jailbreak [142, 143, 144, 145, 146, 147, 148, 149]. In this section, we focus on these three jailbreak methods, especially the multilingual jailbreak. The first two methods involve generic attacks on LLMs, and the latter emphasizes jailbreaking through multiple languages. All these methods are aimed at bypassing the security measures of LLMs to generate malicious information. Since most of the jailbreaking methods are customized, the effectiveness of each method varies depending on the specific LLMs. To provide a comprehensive comparison, we investigate the performance of different jailbreaking methods across various LLMs based on a unified evaluation framework, as shown in Table 4.
@@ 6.1.1 Greedy Coordinate Gradient Jailbreak
The Greedy Coordinate Gradient  [134] is an attack method based on greedy algorithms and gradient metric design hints. It first creates a seed prompt, and then iteratively replaces the tokens. The best prompt for attacking the model is determined by calculating the gradient. Sitawarin et al. [133] propose GCG++ to increase the attack success rate (ASR) of the original GCG method on LLaMA-2-7B. On the basis of the original GCG, they replace the cross entropy loss with multi-class loss to avoid vanishing gradients in the $Softmax$, which is effective in attacking LLM. Another discovery is that LLMs such as LLaMA-2-7B are format sensitive, having a strong prior for predicting a space token at the beginning of the model response. GCG++ adds a simple format-ware target string (i.e., “[ASSISTENT]:” before the input prompt to force the model to output harmful content. Through the two changes, GCG++ increases the GCG attack success rate on LLaMA-2-7B from 56% to 80%.
@@ 6.1.2 Prompt-Based Jailbreak
The prompt-based methods aim to mislead LLM into generating harmful content based on specifically designed templates. Among them, LLM-based [140, 406, 133, 401, 138, 398, 402, 407] and rule-based [135, 399, 136, 137, 144] are the two common used methods for designing attacking prompts. The differences between GCG, LLM-based, and rule-based are presented in Table 5 with various aspects.
Rule-based jailbreak methods are those works that design specific prompts to formulate rules for LLMs [135, 399], or directly collect prompts from various channels such as websites and forums [136, 137, 144]. Rule-based jailbreak methods rely more on human participation because these rules are customized based on empirical experience in specific scenarios. Thus, the produced prompts are effective in a narrow area but lack versatility and are static, which can be easily defended.
LLM-based jailbreak methods leverage the instances of LLMs to optimize existing attack prompts or generate new jailbreak prompts. Different from the GCG series approaches, the principle of LLM-based jailbreak methods acts as using an agent LLM to attack a target LLM. Existing studies  [140, 406, 133, 401] usually take the agent LLM as an optimizer, scoring module or evaluator to optimize prompts for attackers. Some other methods [138, 398, 402, 407] use the generation capabilities of LLM to perform multiple operations such as rewriting and shortening feasible attack prompts. They also wrap harmful prompts with scenarios to avoid safe alignment. These operations based on LLMs can enable the original attack prompt to evade the security alignment policy.
@@ 6.1.3 Multilingual Jailbreak
Due to the unequal security alignment of LLMs across different languages, with higher security in high-resource languages compared to low-resource languages, the multilingual jailbreak methods [142, 143, 145, 146, 147] attend to cheat the LLMs via the alignment vulnerabilities between high-resource and low-resource language. Some other methods [148, 149, 403, 135] use special encoding or ciphers to disguise harmful prompts to achieve jailbreak, which can be considered as a type of special language.
@@
The existing studies on the potential security risks of LLMs, primarily focused on the usage of English scenarios. Under the multilingual scenario, Deng et al. [143] reveal the presence of multilingual jailbreak challenges within LLMs, who divide potential risks into two scenarios: (1) unintentional and (2) intentional. The unintentional scenario involves users inadvertently bypassing the safety alignment by querying LLMs with non-English prompts, while the intentional scenario refers to users intentionally combining malicious instructions with multilingual prompts. Compared to high-resource languages, low-resource languages are approximately three times more likely to encounter harmful content in GPT-3.5 and GPT-4, and thus the LLMs are much easier to attack via low-resource languages. This could also be the reason why the model attack approaches are designed with low-resource languages within these two scenarios.
@@
For the unintentional scenario, Shen et al. [142] indicate that the bottleneck in cross-lingual alignment lies within the training stage of LLMs. Thus, they study the effect of instruction tuning with RLHF and supervised fine-tuning (SFT) on the HH-RLHF dataset [408]. The results show that although training with high-resource languages can improve model alignment, the effect of training with low-resource languages is still negligible. For the intentional scenario, various prompts are designed to bypass the security defenses of the LLMs via high-resource languages. In particular, Yong et al. [145] successfully circumvent the safeguard of GPT-4 by translating unsafe English prompts into low-resource languages. They propose the LRL-Combined Attacks approach to achieve 79% ASR on the AdvBench dataset [134]. Xu et al. [146] study a new type of black box jailbreak attack, Cognitive Overload, which is specifically designed for the cognitive structure and processes of LLMs. With Google Cloud API, they translate original English harmful instructions from AdvBench and MasterKey [138] into 52 other languages and propose multilingual cognitive overload [409, 410] to hinder in-context learning and inference process when the knowledge exceeds the limited capacity of LLMs [411]. By using Google Cloud API to translate English to other 52 languages and nllb-200-distilled-1.3B to translate other non-English responses back to English, they conducted a multilingual version of their attack method. With multilingual Cognitive Overload, their results show that LLMs are more vulnerable to non-English adversarial prompts. The observation indicates that the further a language is from English, the more effective the malicious prompt conveyed is in attacking LLMs. Moreover, Li et al. [147] conduct an extensive empirical study on multilingual jailbreak attacks, which develops a novel semantic preservation algorithm to create multilingual jailbreak prompts. With these prompts, they reveal patterns in multilingual jailbreak attacks and implement fine-tuning mitigation methods for defending against cross-lingual jailbreak attacks.
@@
Password jailbreak. From the perspective of writing scripts, special encoding or password can be considered as a language conversion, i.e., an English prompt can be converted into a password representation by encryption. Though the password representation can also be converted to English or other languages via decryption, the encrypted prompts cannot be understood directly by humans but by the LLMs, and thus could lead to secure alignment fails. For example, Huang et al. [149] propose the generation exploitation attack to manipulate variations of decoding methods (e.g., greedy decoding and sampling-based decoding), which can easily disrupt model alignment because existing alignment procedures are based on default decoding settings. Vulnerabilities are exposed when configurations of decoding change slightly. Yuan et al. [148] propose a novel framework CipherChat to examine the generalizability of safety alignment to non-natural languages, where the user can chat with LLMs through cipher prompts. Furthermore, Lv et al. [403] introduce a hypothesis for the safety mechanism of LLMs: intent security recognition followed by response generation. Following this hypothesis, they propose CodeChameleon to transform queries into decryptable formats with custom Python functions. This approach enables the modification of the original query and achieves state-of-the-art ASR on 7 LLMs. Besides, Wei et al. [135] propose Jailbroken, which obfuscates the queries using base64 to bypass the safety training of LLMs. They observe that LLMs can understand base64 but cannot defend against attacks under base64.
@@
Other jailbreak. Apart from the aforementioned methods, Shayegani et al. [412] combine an image targeted towards toxic embeddings with generic prompts to accomplish the jailbreak, which utilizes four embedding space target strategies to poison the vision encoder. Their attacks achieve 87% and 63.3% ASR on two vision-language models, LLaVA [325] and LLaMA-Adapter V2 [413], respectively. Rando et al. [414] embed a “jailbreak backdoor” into the LLM by poisoning the RLHF training data. Then, users can easily achieve jailbreak by using a trigger word like the “sudo” command in the Linux system. Wolf et al. [415] propose a theoretical approach Behavior Expectation Bounds (BEB) that investigates several limitations of alignment in LLMs, which exposes fundamental problems and emphasizes the necessity of designing reliable mechanisms to ensure AI safety.
Model Attacker: PAIR [401]
No Defense SmoothLLM [150] Perplexity Filter [152]
@@
Table 6: An overview of defense methods under jailbreaking on listed LLMs. The evaluation method for the Attack Success Rate (ASR) indicator is consistent with JailbreakBench [417]. The responses of LLMs are evaluated using LLaMAGuard-7B.